A body is thrown vertically upward and attains a velocity of 15m/s at the half the maximum height, the maximum height up to which the body can reach will be
A. 45.8m
B. 34.5m
C. 22.5m
D. 17.8m

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Sqr(v ) = sqr(u)- 2(gs)

S= (15*15/20)

S= h/2

h = 22.5 m

Regards